(1/4) + (1/8)= 3/8
We generally do it by taking l.c.m of 4& 8 and do further by general procedure.
Can we do it by multipling 4 and 8 and then divide the result 32 by 4 and by 8. We get again 3/8. Is this method correct?
1
Expert's answer
2020-05-06T18:57:59-0400

Yes it's correct, because

"\frac{1}{4}+\frac{1}{8}=\frac{8+4}{4*8}=\frac{12}{32}=\frac{3*4}{8*4}=\frac{3}{8}*\frac{4}{4}=\frac{3}{8}*1=\frac{3}{8}"